# Counterfactual Reasoning for Multi-Label Image Classification via Patching-Based Training

### v1

@@ 6.5 Step-Wise Evaluation
We perform a step-wise evaluation on COCO to have a better understanding of how TDE keep the positive side while removing the negative side of the mediated effect. Figure 6 (a) shows there is no significant difference between the performance of original logits $\bm{p}$ and patch logits $\bm{q}$. In some classes, the former performs better, while in other classes, the latter exhibits superior performance. This means that introducing the patching technique does not lead to an improvement in model performance. Figure 6 (b) illustrates the improvements in performance of the proposed TDE logits on the basis of the optimum between the original logits $\bm{p}$ and patch logits $\bm{q}$. We can see TDE maintains positive effects in almost every class, achieving the retention of positive causal effect and the removal of negative one.
Figure 5: Conditional FPR and conditional TPR of the label pairs with co-occurrence probabilities larger than 0.2 on COCO. The pair indices are sorted according to the performance of Baseline. PAT-T outperforms Baseline in around 76% and 77% percentage of label pairs in terms of CFPR and CTPR.
@@ 6.6 Case Studies
(a) Original Probability $\bm{p}$ vs Patch Probability $\bm{q}$
(b) Optimum between $(\bm{p},\bm{q})$ vs TDE Probability $\bm{p}+\lambda\cdot\bm{q}$
Figure 6: Corresponding to three output logits, original logits $\bm{p}$ on an image, patch logits $\bm{q}$ on patches, and TDE logits $\bm{p}+\bm{q}$, we illustrate the per-class average precision increments between performance of each two logits to show the effectiveness of TDE.
@@ 7 Conclusion
The paper studies the problem of multi-label learning, which aims to improve model performance by leveraging label correlations. Previous studies demonstrate that the powerful fitting capacity of DNNs often leads to overfitting to co-occurrence relationships. We establish a causal reasoning framework that treats the correlative features caused by the target object and its co-occurring ones as a mediator, exhibiting both positive and negative impact on model predictions. To keep the positive side while removing the negative impact, we propose to enhance the direct causal effect caused by the target object when masking its co-occurring objects. In implementation, we propose patching-based training to identify the pivot patch that contains the target object. Experimental results on multiple benchmark datasets validate the superiority of our method to state-of-the-art methods.
@@ Appendix C Pseudocode of PAT-I and PAT-T
Algorithm 1 and Algorithm 2 provide the pseudocode of PAT-I and PAT-T in a Pytorch-like style, respectively. We can see that the algorithms are simple to implement.

### v2

@@ 6.5 Step-Wise Evaluation
We perform a step-wise evaluation on MS-COCO to have a better understanding of how TDE keeps the positive side while removing the negative side of the mediated effect. Figure 6 (a) shows there is no significant difference between the performance of original logits $\bm{p}$ and patch logits $\bm{q}$. In some classes, the former performs better, while in other classes, the latter exhibits superior performance. This means that introducing the patching technique does not lead to an improvement in model performance. Figure 6 (b) illustrates the improvements in performance of the proposed TDE logits on the basis of the optimum between the original logits $\bm{p}$ and patch logits $\bm{q}$. We can see TDE maintains positive effects in almost every class, achieving the retention of positive causal effect and the removal of negative one.
(a) Original Prediction $\bm{p}$ vs Patch Prediction $\bm{q}$
(b) Optimum between $(\bm{p},\bm{q})$ vs TDE Prediction $\bm{p}+\lambda\cdot\bm{q}$
Figure 6: Corresponding to three output logits, original logits $\bm{p}$ on an image, patch logits $\bm{q}$ on patches, and TDE logits $\bm{p}+\lambda\cdot\bm{q}$, we illustrate the per-class average precision increments between performance of each two logits to show the effectiveness of TDE.
@@ 6.6 Case Studies
Table 4: Comparison results on MS-COCO between patch logits $\bm{q}$ and TDE logits $\bm{p}+\lambda\cdot\bm{q}$.
Backbone Patch Logits TDE Logits
TResNetL 87.1 88.7
TResNetL (21k) 88.8 89.2
TResNetL (OI) 89.5 90.8
Q2L-TResL 87.9 89.5
Q2L-TResL (21k) 89.1 91.0
Q2L-TResL (OI) 90.1 91.7
6.7 Study on Influence of Image Scale
In the PAT method, we found that training a backbone with images and patches results in a significant decline in performance. This is due to their inconsistent sizes, i.e., the patches are only a quarter of the size of the image, severely disrupting feature representation learning. To deal with this problem, we first resize the patches to the size of the original image before inputting them into the model. Note that this operation does not incorporate additional information, because we first crop the resized image into patches and then resize them again. This operation only affects the scale of the targets. We perform experiments to verify that the main reason for the performance improvement is not due to the increase in object scale. Specifically, we train a model based on only resized patches using the loss function $\ell(\delta(\bm{q}),\bm{y})$ in Eq. (5). Table 4 reports mAP scores of patch logits and TDE logits. From the table, we can see that when the representation capability of the backbone is relatively weak, patch logits performs slightly better than the image logits (baseline results in Table 1 with corresponding backbones). As the representation capability of the backbone increases, this advantage disappears, and there is no significant difference in performance between patch logits and image logits. While our TDE logits still outperform patch logits with a large margin. These results convincingly validate that the effectiveness of our method stems from decoupling co-occurrence relationships using TDE.
@@ 7 Conclusion
The paper studies the problem of multi-label learning, which aims to improve model performance by leveraging label correlations. Previous studies demonstrate that the powerful fitting capacity of DNNs often leads to overfitting to co-occurrence relationships. We establish a causal reasoning framework that treats the correlative features caused by the target object and its co-occurring ones as a mediator, exhibiting both positive and negative impacts on model predictions. To keep the positive side while removing the negative impact, we propose to enhance the direct causal effect caused by the target object when masking its co-occurring objects. In implementation, we propose patching-based training to identify the pivot patch that contains the target object. Experimental results on multiple benchmark datasets validate the superiority of our method to state-of-the-art methods.
Impact Statement
This paper presents work whose goal is to advance the field of Machine Learning. There are many potential societal consequences of our work, none which we feel must be specifically highlighted here.
Acknowledgements
Sheng-Jun Huang was supported by Natural Science Foundation of Jiangsu Province of China (BK20222012, BK20211517), the National Key R&D Program of China (2020AAA0107000), and NSFC (62222605). Masashi Sugiyama was supported by JST CREST Grant Number JPMJCR18A2.
@@ Acknowledgements
Algorithm 1 Pseudocode of PAT-I in a PyTorch-like style
  # f: model
  # t: temperature
  # lam: tde parameter
  for [x; u_1; u_2; u_3; u_4], y in loader:
     # image logits and patch logits NxCx5
     [l_x; l_u_1; l_u_2; l_u_3; l_u_4] = f([x; u_1; u_2; u_3; u_4])
     # patch weights NxCx4
     [w_1; w_2; w_3; w_4]=softmax([l_u_1;l_u_2;l_u_3;l_u_4]/t, dim=-1)
     # tde logits
     l_tde = l_x + lam*(w_1*l_u_1 + w_2*l_u_2 + w_3*l_u_3 + w_4*l_u_4).sum(dim=-1)
     # mAP score calculation
     mAP_score = mAP_function(l_tde, y)
Algorithm 2 Pseudocode of PAT-T in a PyTorch-like style
  # f: model
  # t: temperature
  for [x; u_1; u_2; u_3; u_4], y in loader:
     # image logits and patch logits NxCx5
     [l_x; l_u_1; l_u_2; l_u_3; l_u_4] = f([x; u_1; u_2; u_3; u_4])
     # patch weights NxCx4
     [w_1; w_2; w_3; w_4]=softmax([l_u_1;l_u_2;l_u_3;l_u_4]/t, dim=-1)
     # weighted logits
     l_w = (w_1*l_u_1 + w_2*l_u_2 + w_3*l_u_3 + w_4*l_u_4).sum(dim=-1)
     # loss, Eq.(5)
     loss = AsymmetricLoss(l_x, y) + AsymmetricLoss(l_w, y)
     # model update
     loss.backward()
@@ Appendix C Pseudocode of PAT-I and PAT-T
Algorithm 1 and Algorithm 2 provide the pseudocode of PAT-I and PAT-T in a Pytorch-like style, respectively. We can see that both two algorithms are simple to implement and act like plug-and-play plugins, which bring clear benefits.